# Federated Continual Learning Goes Online: Leveraging Uncertainty for Modality-Agnostic Class-Incremental Learning

### v2

@@ 4.2 Empirical Results
LC A()($\uparrow$)( ↑ ) 47.66 $\pm$ 2.52 47.24 $\pm$ 4.48 56.49 $\pm$ 2.21 51.68 $\pm$ 2.18
F()($\downarrow$)( ↓ ) 41.42 $\pm$ 3.64 36.51 $\pm$ 6.86 26.65 $\pm$ 6.30 33.94 $\pm$ 8.80
MS A()($\uparrow$)( ↑ ) 47.87 $\pm$ 1.95 49.35 $\pm$ 2.57 52.82 $\pm$ 3.51 51.98 $\pm$ 2.42
F()($\downarrow$)( ↓ ) 36.96 $\pm$ 4.74 41.18 $\pm$ 7.67 30.42 $\pm$ 4.05 33.07 $\pm$ 2.59
RC A()($\uparrow$)( ↑ ) 50.42 $\pm$ 3.92 48.96 $\pm$ 2.13 55.93 $\pm$ 5.85 53.23 $\pm$ 1.58
F()($\downarrow$)( ↓ ) 32.95 $\pm$ 4.04 39.84 $\pm$ 5.70 26.57 $\pm$ 6.35 29.90 $\pm$ 7.36
EN A()($\uparrow$)( ↑ ) 49.18 $\pm$ 2.36 48.58 $\pm$ 2.09 53.80 $\pm$ 4.56 52.80 $\pm$ 2.77
F()($\downarrow$)( ↓ ) 29.55 $\pm$ 6.39 36.78 $\pm$ 5.75 29.48 $\pm$ 12.65 35.76 $\pm$ 10.16
BI A()($\uparrow$)( ↑ ) 49.18 $\pm$ 2.61 59.72 $\pm$ 2.39 50.16 $\pm$ 2.79 59.06 $\pm$ 2.55
F()($\downarrow$)( ↓ ) 36.01 $\pm$ 7.81 18.84 $\pm$ 8.64 37.28 $\pm$ 9.50 23.10$\pm$ 9.81
Table 3: Comparison of average last accuracy (A) and last forgetting (F) on KC-Cell (4 tasks).
@@
Table 4: Comparison of average last accuracy (A) and last forgetting (F) on 20NewsGroups (5 tasks).
ER CBR LC MS RC EN BI
M=60 A()($\uparrow$)( ↑ ) 42.33 $\pm$ 1.66 45.21 $\pm$ 1.86 42.90 $\pm$ 1.23 44.14 $\pm$ 1.63 44.15 $\pm$ 1.32 41.65 $\pm$ 1.82 44.92 $\pm$ 1.63
@@
5 Conclusion and Limitations
In this work, we investigate the challenges of real-world federated continual learning contexts, specifically when data arrive in streams of small chunks and the local models need to be updated with high frequency. Current research in FCL focuses on developing generative-based solutions that imply an offline setting where generators are trained at the end of each task. However, we advocate that in realistic conditions, to keep the model up-to-date and reduce the communication overhead, local models should be trained every time new data are received. For this, we devise and formalise a new scenario for the online problem in FCL. To solve catastrophic forgetting, we introduce a simple memory-based baseline that, via a combination of uncertainty-based update and random replay, outperforms state-of-the-art approaches. The newly proposed BI-based for epistemic uncertainty shows its superiority in reducing CF compared to standard uncertainty metrics, it is simple to implement and applicable across a wide range of data modalities. Furthermore, we validate our findings on more probing scenarios from the biomedical domain. This confirms its ability to sample more robust and representative samples in challenging tasks.
@@ 5 Conclusion and Limitations
A limitation of the proposed approach is the need to use an ensembling technique (in our case, TTA) in order to estimate the BI scores. This may result in a reduced efficiency compared to, e.g., ER. Furthermore, the TTA-based estimation of the Bregman Information is only an estimate of the true unknown uncertainty; the estimator used throughout the experiments (eq. 1) is only asymptotically unbiased, and underestimates the theoretical quantity (Gruber and Buettner, 2023).
@@
Appendix A Appendix / supplemental material
@@ A.3 Dataset Statistics
In Table 5, we report the statistics of the datasets used for the main experiments. We use dataset from different domains (CIFAR10, CRC-Tissue, KC-Cell) and with different modalities (20NewsGroups). For CRC-Tissue, in order to create tasks with decreasing size and equal number of classes, we remove the smallest class from our evaluation.

### v3

@@ 4.2 Empirical Results
LC A()($\uparrow$)( ↑ ) 47.66 $\pm$ 2.52 56.02 $\pm$ 4.89 56.49 $\pm$ 2.21 58.12 $\pm$ 2.94
F()($\downarrow$)( ↓ ) 41.42 $\pm$ 3.64 13.25 $\pm$ 4.91 26.65 $\pm$ 6.30 11.13 $\pm$ 3.12
MS A()($\uparrow$)( ↑ ) 47.87 $\pm$ 1.95 54.92 $\pm$ 2.93 52.82 $\pm$ 3.51 55.64 $\pm$ 4.42
F()($\downarrow$)( ↓ ) 36.96 $\pm$ 4.74 17.24 $\pm$ 8.46 30.42 $\pm$ 4.05 9.59 $\pm$ 6.93
RC A()($\uparrow$)( ↑ ) 50.42 $\pm$ 3.92 57.50 $\pm$ 3.01 55.93 $\pm$ 5.85 58.81 $\pm$ 3.34
F()($\downarrow$)( ↓ ) 32.95 $\pm$ 4.04 12.91 $\pm$ 5.00 26.57 $\pm$ 6.35 8.71 $\pm$ 5.86
EN A()($\uparrow$)( ↑ ) 49.18 $\pm$ 2.36 55.22 $\pm$ 4.05 53.80 $\pm$ 4.56 58.48 $\pm$ 3.06
F()($\downarrow$)( ↓ ) 29.55 $\pm$ 6.39 14.95 $\pm$ 8.85 29.48 $\pm$ 12.65 8.16 $\pm$ 3.79
BI A()($\uparrow$)( ↑ ) 49.18 $\pm$ 2.61 57.05 $\pm$ 3.65 50.16 $\pm$ 2.79 62.33 $\pm$ 1.93
F()($\downarrow$)( ↓ ) 36.01 $\pm$ 7.81 13.63 $\pm$ 8.04 37.28 $\pm$ 9.50 7.99 $\pm$ 4.08
KC-Cell
@@
Table 4: Comparison of average last accuracy (A) and last forgetting (F) on text classification tasks.
Dataset ER CBR LC MS RC EN BI
20NewsGroup M=60 A()($\uparrow$)( ↑ ) 42.33 $\pm$ 1.66 45.21 $\pm$ 1.86 42.90 $\pm$ 1.23 44.14 $\pm$ 1.63 44.15 $\pm$ 1.32 41.65 $\pm$ 1.82 44.92 $\pm$ 1.63
@@
DBPedia M=60 A()($\uparrow$)( ↑ ) 75.99 $\pm$ 2.10 76.24 $\pm$ 1.21 74.73 $\pm$ 2.04 75.21 $\pm$ 1.92 75.36 $\pm$ 1.61 74.58 $\pm$ 1.88 77.86 $\pm$ 1.95
F()($\downarrow$)( ↓ ) 24.73 $\pm$ 2.70 23.06 $\pm$ 1.58 27.06 $\pm$ 2.30 26.44 $\pm$ 2.28 26.52 $\pm$ 1.77 27.14 $\pm$ 2.15 23.06 $\pm$ 2.10
M=100 A()($\uparrow$)( ↑ ) 76.78 $\pm$ 1.04 79.78 $\pm$ 0.83 76.21 $\pm$ 1.58 76.18 $\pm$ 1.70 76.18 $\pm$ 1.50 76.33 $\pm$ 1.45 78.68 $\pm$ 1.08
F()($\downarrow$)( ↓ ) 24.14 $\pm$ 1.18 21.45 $\pm$ 1.49 25.15 $\pm$ 1.60 25.19 $\pm$ 1.83 25.29 $\pm$ 1.78 24.87 $\pm$ 1.83 21.79 $\pm$ 1.41
Yahoo Answers M=500 A()($\uparrow$)( ↑ ) 45.80 $\pm$ 1.77 47.45 $\pm$ 2.86 47.87 $\pm$ 3.38 47.60 $\pm$ 3.38 48.25 $\pm$ 2.61 49.32 $\pm$ 3.61 50.13 $\pm$ 1.69
F()($\downarrow$)( ↓ ) 40.15 $\pm$ 2.91 36.38 $\pm$ 4.29 36.02 $\pm$ 6.81 35.69 $\pm$ 4.61 34.60 $\pm$ 3.49 33.67 $\pm$ 5.12 32.65 $\pm$ 3.21
M=1000 A()($\uparrow$)( ↑ ) 45.48 $\pm$ 2.11 44.62 $\pm$ 3.23 46.40 $\pm$ 3.07 47.00 $\pm$ 2.75 46.33 $\pm$ 3.68 46.70 $\pm$ 1.11 47.03 $\pm$ 3.36
F()($\downarrow$)( ↓ ) 39.40 $\pm$ 3.47 39.17 $\pm$ 4.16 37.15 $\pm$ 5.78 37.15 $\pm$ 4.35 38.08 $\pm$ 5.26 37.85 $\pm$ 1.78 37.25 $\pm$ 4.17
Table 5: Comparison of average last accuracy (A) and last forgetting (F) on vision and textual tasks for standard FL, generative-based FCL, and our method.
Method CIFAR10 CIFAR100 CRC-Tis. KC-Cell 20News DBPedia Yahoo
FedAvg A()($\uparrow$)( ↑ ) 16.90 $\pm$ 0.55 3.57 $\pm$ 0.33 22.26 $\pm$ 2.29 15.04 $\pm$ 0.83 6.34 $\pm$ 0.58 19.57 $\pm$ 0.30 14.18 $\pm$ 1.59
F()($\downarrow$)( ↓ ) 78.29 $\pm$ 2.60 19.73 $\pm$ 0.99 87.39 $\pm$ 4.06 74.08 $\pm$ 3.70 58.84 $\pm$ 1.91 99.02 $\pm$ 0.32 70.90 $\pm$ 2.88
Weighted FedAvg A()($\uparrow$)( ↑ ) 16.51 $\pm$ 0.77 4.22 $\pm$ 0.20 22.00 $\pm$ 2.40 14.85 $\pm$ 0.92 6.36 $\pm$ 0.56 19.23 $\pm$ 0.44 12.52 $\pm$ 1.60
F()($\downarrow$)( ↓ ) 70.08 $\pm$ 1.57 20.18 $\pm$ 0.23 80.40 $\pm$ 5.15 76.21 $\pm$ 2.72 50.22 $\pm$ 2.41 97.64 $\pm$ 1.25 69.00 $\pm$ 2.11
FedProx A()($\uparrow$)( ↑ ) 16.57 $\pm$ 0.50 3.73 $\pm$ 0.15 22.12 $\pm$ 1.26 14.79 $\pm$ 1.01 12.25 $\pm$ 0.40 19.58 $\pm$ 0.26 17.35$\pm$ 1.34
F()($\downarrow$)( ↓ ) 76.74 $\pm$ 1.12 20.26 $\pm$ 0.58 87.86 $\pm$ 4.78 73.38 $\pm$ 3.66 67.31 $\pm$ 3.14 99.06 $\pm$ 0.32 82.54 $\pm$ 0.55
MFCL A()($\uparrow$)( ↑ ) 16.42 $\pm$ 0.32 4.60 $\pm$ 0.09 27.62 $\pm$ 7.99 16.32 $\pm$ 1.21
F()($\downarrow$)( ↓ ) 55.64 $\pm$ 3.03 13.69 $\pm$ 0.33 76.34 $\pm$ 12.68 70.88 $\pm$ 1.94
FedCIL A()($\uparrow$)( ↑ ) 16.26 $\pm$ 1.16 2.45 $\pm$ 0.28 22.44 $\pm$ 1.23 13.14 $\pm$ 1.31
F()($\downarrow$)( ↓ ) 49.87 $\pm$ 0.61 4.99 $\pm$ 0.81 62.47 $\pm$ 1.35 55.82$\pm$ 1.60
BI (best) A()($\uparrow$)( ↑ ) 35.83 $\pm$ 2.60 14.31 $\pm$ 0.76 62.33 $\pm$ 1.93 21.61 $\pm$ 1.23 46.72 $\pm$ 2.11 78.68 $\pm$ 1.08 50.13 $\pm$ 1.69
F()($\downarrow$)( ↓ ) 19.07 $\pm$ 2.17 6.96 $\pm$ 1.40 7.99 $\pm$ 4.08 59.05 $\pm$ 4.42 26.97 $\pm$ 2.36 21.79 $\pm$ 1.41 32.65 $\pm$ 3.21
5 Discussion and Conclusion
The novelty of the online scenario for FCL tasks raises questions on different aspects of the proposed approach and we next provide an in-depth discussion of key methodological choices and findings.
Ablation studies - Appendix A.5. We conduct several ablation studies to assess the impact of different hyperparameters (i.e., burn-in period, jump parameter, parameter averaging strategy, and augmentation set) on the learning performance of the proposed approach while also providing insights and explanations for usage best practices in this newly introduced context. One key aspect is the set of augmentations. From the results, the number of augmentations is important to improve the uncertainty estimation and, consequently, the performance of the approach. In terms of the choice of modality-specific augmentations, we rely on the large body of literature on TTA and show that standard augmentations work well both in the medical domain and for standard benchmarks of natural images. For other modalities, we propose to perform the augmentation on the latent representation level by simply adding Gaussian noise.
Practicality and computational time - Appendices A.6 and A.7. In terms of practicality, for standard datasets like CIFAR10 and CIFAR100, the number of communication rounds per task with a jump parameter $q=5$ is comparable to the one reported in FedCIL (40 communication rounds per task) and considerably smaller than MFCL (100 communication rounds per task). The same consideration is also valid for the training computational time; our approach has a comparable computational time compared to FedCIL and a considerably smaller one compared to MFCL.
Performance of generative-based approaches - Appendix A.8. The performance gap of generative-based solutions in our experiments stems from the different assumptions of the online scenario compared to the original setting. For instance, MFCL on CIFAR100 assumes 100 communication rounds for each task. This implies that the whole task dataset is processed 100 times, allowing an effective synthetic image generation. In our online scenario, instead, we have access to one batch of 10 images at every iteration. In this case, the effectiveness of generative-based solutions is limited by their need for large amounts of data. This is in line with the online-CL literature where the most successful solutions rely on memory-based approaches (Ma et al., 2022; Soutif-Cormerais et al., 2023). To illustrate this point, we generate some images using MFCL trained in the online setting (see Figure 5 in Appendix A.8). The synthetic images provide limited information about past tasks resulting in MFCL being better than standard FL approaches without memory, but not competitive compared to our solution (see Table 5).
@@ 5 Discussion and Conclusion
Limitations. A limitation of the proposed approach is the need to use an ensembling technique (in our case, test-time augmentation) in order to compute the BI-based estimates. This may result in a reduced efficiency compared to, e.g., ER. Furthermore, the TTA-based estimation of the Bregman Information is only an estimate of the true unknown uncertainty; the estimator used throughout the experiments (Eq. 1) is only asymptotically unbiased and underestimates the theoretical quantity (Gruber and Buettner, 2023).
Conclusions. In this work, we address the challenges of real-world FCL, where new data arrive in streams of small chunks that cannot be fully retained after being processed during training. While current research in FCL focuses on generative-based solutions that imply an offline setting where generators are trained at the end of each task, we advocate that in realistic conditions complete task datasets may be unavailable and local models should be updated every time new data are received. For this, we devise and formalise a new scenario for the online problem in FCL. To solve CF, we introduce an effective memory-based baseline that combines uncertainty-aware updates with random replay, outperforming state-of-the-art methods. Unlike generative solutions, the newly proposed BI-based estimation for epistemic uncertainty is simple to implement and applicable across different data modalities. Furthermore, compared to standard uncertainty metrics, the empirical results show its superiority in reducing CF in standard settings as well as imbalanced and probing real-world scenarios taken from the biomedical domain. This confirms the ability of the proposed uncertainty-aware strategy to sample more robust and representative samples in challenging tasks and opens new frontiers in FCL, particularly for domains where data scarcity and imbalance are prevalent.
@@
Appendix A Appendix
@@ A.3 Dataset Statistics
In Table 6, we report the statistics of the datasets used for the main experiments. We use dataset from different domains (CIFAR10, CIFAR100, CRC-Tissue, KC-Cell) and with different modalities (20NewsGroups, Yahoo Answers, DBPedia). For CRC-Tissue, in order to create tasks with decreasing size and equal number of classes, we remove the smallest class from our evaluation. For DBPedia and Yahoo Answers Zhang et al. [2015], we randomly select a subset of 25000 and 33600 samples respectively.
@@ A.4 Evaluation Metrics
A.5 Ablation Study
In this section, we investigate the effect of various hyperparameters (i.e., burn-in period, jump parameter, parameter averaging strategy, and set of augmentations) on the learning capabilities of the proposed approach while also providing insights and explanations for best practices to use them in this newly proposed context.
Burn-in period. From the results reported in Table 7, the burn-in period seems to provide some benefit compared to the standard case (burn-in equal to 1) although its contribution is not crucial for improving the overall performance of the approach.
Table 7: Effect of the burn-in period on accuracy (A) and forgetting (F).
Burn-in A($\uparrow$) F($\downarrow$)
1 23.77 29.05
10 25.64 29.64
30 24.89 35.77
50 22.43 34.71
100 24.48 33.20
Parameter averaging strategy. The parameter averaging strategy can be changed without affecting particularly the performance of the approach as can be seen in Table 8. Overall, taking the average between the current and the previous global parameters (as proposed in Shenaj et al. [2023]) helps in reducing CF compared to the vanilla FedAvg.
Table 8: Effect of the parameter averaging strategy on accuracy (A) and forgetting (F).
Strategy A($\uparrow$) F($\downarrow$)
FedAvg 26.73 44.38
Weighted FedAvg 24.89 37.77
FedProx 26.67 38.32
Jump parameter. A value between 3 and 5 is a good trade-off for communication efficiency and good predictive performance. If the communication frequency is troublesome due to hardware constraints, one might use higher values at the price of losing pure predictive performance.
Table 9: Effect of the jump parameter $q$ on accuracy (A) and forgetting (F).
$q$ A($\uparrow$) F($\downarrow$)
1 21.15 16.55
3 37.83 34.84
5 24.89 35.77
10 22.75 46.46
50 21.99 37.86
Set of augmentations. To ablate the influence of the augmentations on the performance of our approach, starting from the set of augmentations listed in Appendix A.2, we randomly remove 4 and 2 augmentations from the list. The results reported in Table 10 suggest that removing augmentations decreases the performance of the proposed approach. This is probably because the higher is the number of augmentations used, the more precise is the uncertainty estimation resulting in better samples stored in the memory.
Table 10: Effect of the augmentation set on accuracy (A) and forgetting (F).
Nr. of augmentations A($\uparrow$) F($\downarrow$)
8 20.84 40.96
10 22.18 38.03
12 24.89 35.77
A.6 Practicality Analysis - Number of Communication Rounds
For standard datasets like CIFAR10 and CIFAR100, the number of communication rounds per task with a jump parameter $q=5$ (as used in the main experiments) is comparable to the one proposed in FedCIL (40 communication rounds per task) and considerably smaller than MFCL (100 communication rounds per task). More in details, for each of the aforementioned datasets, each client possesses roughly 2000 images. With a batch size equal to 10, this results in 200 batches per task. Thus, removing the burn-in period of 30 batches, we are left with 170 batches. This means that, for a given task, the number of communication rounds with $q=5$ is approximately 35. The frequency can be further reduced by increasing the jump parameter with a slight degradation in the predictive performance of the model (see results in the ablation study reported in Appendix A.5 Table 9.).
Table 11: Communication rounds for FCL approaches on CIFAR10 and CIFAR100.
Communication rounds
MFCL 100
FedCIL 40
O-FCL (ours) 35
A.7 Time Complexity Analysis
Table 12 reports the runtime in seconds at the end of the training procedure on CIFAR10 for FedAvg (no replay mechanism), ER (random, memory-based), CBR (class-balanced, memory-based), MFCL (generative-based), FedCIL (generative-based), and BI (uncertainty-based, memory-based). From the results, the computational time of BI is in line with most of other uncertainty-based approaches and much shorter than generative approaches. The increase in runtime from CBR to BI stems from the requirement to generate TTA images. In comparison to generative approaches, the runtime of BI is similar or, compared to MFCL, considerably shorter.
Table 12: Runtime (in seconds) at the end of the learning procedure on CIFAR10.
Runtime (in seconds)
FedAvg $\sim$ 190
ER $\sim$ 210
CBR $\sim$ 300
MFCL $\sim$ 3840
FedCIL $\sim$ 600
BI (ours) $\sim$ 660
A.8 Performance Analysis of Generative-based Models
The proposed online-FCL (O-FCL) scenario is very different from the standard FCL scenario presented in, e.g., MFCL [Babakniya et al., 2024] and related works based on generative-based solutions. In the standard scenario, they assume to work on task datasets that can be iterated multiple times. For instance, let us consider MFCL on CIFAR100 in which the authors assume to have 100 communication rounds for each task. This implies that the whole task dataset is processed 100 times. Thus, after many iterations over the same task dataset, the generative model can effectively generate synthetic images for the given task. In our online scenario, instead, we only have access to one batch of 10 images at every iteration. In this setting, generative-based solutions are limited in their capabilities as they require many data points and many iterations over the training data for being effective. For this reason, we believe that the performance gap compared to the standard FCL setting is given by the inefficiency of generative models in online settings where we do not have the possibility to retrain the model on the same data multiple times. This is in line with the literature in online-CL where the most successful solutions rely on memory-based solutions [Ma et al., 2022, Soutif-Cormerais et al., 2023]. To prove this point, we generated some images using MFCL at the end of the training procedure in the online setting (see Figure 5). Differently from the images reported in the original paper, we can see that no informative pattern has been discovered; the synthetic images provide limited information about the past information. This is also corroborated by the results of MFCL on the datasets for vision tasks reported in the main paper. From the results, we can see that MFCL is better than standard approaches without memory, but it is not competitive compared to memory-based approaches.
Figure 5: Synthetic images generated by MFCL [Babakniya et al., 2024] in the online setting.
A.9 Pseudocode of O-FCL
Algorithm 1 O-FCL learning scheme
1:Input K: number of clients, T: number of tasks, $q$: jump parameter, burn-in parameter, $bs$: batch size.
2:Notation $k$: current client; $\mathcal{M}_{k}$: client memory, $\left[\bm{C}\right]$: client set, $t$: current task, $b_{k}^{t}$ current batch size for client $k$ and task $t$, $\bm{m}$: memory sample (size equal to $bs$); $bn_{k}^{t}$: number of batches for client $k$ and task $t$, $\bm{\theta}_{g}$: global model parameters, $\bm{\theta}_{k}$: local model parameters.
3:while training not completed for all clients do
4:
5:     Client training
6:     for $k\in[1,\dots,K]$ do $\triangleright$ For each client
7:         $b_{k}^{t},bn_{k}^{t}\leftarrow$ GetNextBatch($\bm{C}_{k}$) $\triangleright$ Get one batch from $t$
8:         if $t=0$ then
9:              $\bm{\theta}_{k}\leftarrow$ Train($b_{k}^{t}$)
10:         else
11:              $\bm{m}\leftarrow$ SampleFromMemory($\mathcal{M}_{k}$) $\triangleright$ Random sampling from memory
12:              $\bm{\theta}_{k}\leftarrow$ Train($b_{k}^{t}$, $\bm{m}$)
13:         end if
14:         $\mathcal{M}_{k}\leftarrow$ UpdateMemory($\mathcal{M}_{k}$, $b_{k}^{t}$)
15:     end for
16:
17:     Parameter Averaging
18:     if $bn_{k}^{t}>$ burn-in and $bn_{k}^{t}\ \%\ q=0$ then
19:         $\bm{\theta}_{g}\leftarrow$ AggrModel($\bm{\theta}_{k}\quad\forall k\in[1,\dots,K]$)
20:         for $k\in[1,\dots,K]$ do
21:              $\bm{\theta}_{k}\leftarrow$ UpdateParameters($\bm{\theta}_{g}$)
22:         end for
23:     end if
24:end while
25:
26:function UpdateMemory($\mathcal{M}_{k}$, $b_{k}^{t}$)
27:     for $c\in b_{k}^{t}$ do $\triangleright$ For each label $c$ in the batch size
28:         $\bm{m}_{c}\leftarrow$ $\mathcal{M}_{k}[y=c]$triangleright Extract samples with label $c$ from memory
29:         $\bm{a}_{c}\leftarrow$ $\bm{m}_{c}\cup b_{k}^{t}[y=c]$triangleright Get candidate samples for label $c$
30:         $\bm{u}_{c}\leftarrow$ $u_{BI}(\textsc{TTA}(\bm{a}_{c}))$triangleright Compute uncertainty with Eq. (1)
31:         $\mathcal{M}_{k}\leftarrow$ ReplaceSamples($\bm{u}_{c}$) $\triangleright$ Update $\mathcal{M}_{k}$ based on $\bm{u}_{c}$
32:     end for
33:     return updated $\mathcal{M}_{k}$
34:end function